# Correlation between $\mathrm{H}\rm{\alpha}$ emitters and their cosmic web environment at $z\sim 1$

### v1

@@ Abstract
Future near-infrared spectroscopic galaxy surveys will target high-redshift emission-line galaxies (ELGs) to test cosmological models. Deriving optimal constraints from emission-line galaxy clustering hinges on a robust understanding of their environmental dependence. Using the TNG300-1 simulation, we explore the correlation between properties of $\mathrm{H}\rm{\alpha}$ emitters and their environment anisotropy rather than traditional density-based measures. Our galactic $\mathrm{H}\rm{\alpha}$ emission model includes contributions from the warm interstellar medium. The environment anisotropy and type are assigned using a halo mass-dependent smoothing scale. We find that most luminous ELGs ($L_{\mathrm{H}\rm{\alpha}}>10^{42}\ {\rm erg\,s^{-1}}$) reside in filaments and knots. More generally, ELGs are more biased in strongly anisotropic environments. While correlations with galactic properties are found to be weak, they are statistically significant for host halo masses $M\lesssim 10^{12}\ M_{\odot}/h$. Our analysis motivates further investigation into how environmental anisotropy influences galaxy evolution, and highlights the potential for leveraging these effects in the analyses of upcoming cosmological surveys.
1 Introduction
The next generation of galaxy surveys, including Euclid, DESI and SPHEREx (Euclid Collaboration et al., 2024a; Aghamousa et al., 2016; Doré et al., 2014) will provide detailed measurements of emission-line galaxy (ELG) clustering. These data will open new avenues for probing key cosmological scenarios, including the nature of dark matter and dark energy, the physics of the early Universe, and the mass of neutrinos. Interpreting the clustering and physical characteristics of emission line galaxies (ELGs) requires a nuanced understanding of the environments in which they reside. Accurate modeling of ELG clustering further hinges on robust priors for galaxy bias parameters, as recent work has shown that the bias of ELGs is sensitive to both selection effects and the details of galaxy formation physics, motivating the use of simulation-calibrated priors and perturbative approaches (e.g., Merson et al., 2019; Barreira et al., 2020; Zhai et al., 2021; Barreira et al., 2021; Yuan et al., 2022; Marinucci et al., 2023; Euclid Collaboration et al., 2024b; Pei et al., 2024; Madar et al., 2024; Reyes-Peraza et al., 2024; Akitsu, 2024; Ivanov et al., 2024; Garcia-Quintero et al., 2024; Khederlarian et al., 2024; Ortega-Martinez et al., 2025; Ivanov, 2025; Sullivan et al., 2025, for recent studies).
@@ 1 Introduction
Since there is no unique way to define the environment of galaxies, several methods have been developed to characterize it. Common approaches include local density estimators, such as counts-in-cylinders or adaptive smoothing techniques (Dressler, 1980; Cooper et al., 2006) as well as classifications based on group membership or nearest-neighbor distances (e.g., Baldry et al., 2006; Yang et al., 2009). Beyond simple density-based measures (e.g. Hoyle et al., 2005; Park et al., 2007; Blanton & Moustakas, 2009; Tempel et al., 2011), the large-scale tidal field provides a comprehensive framework for characterizing the cosmic web (Shen et al., 2006; Hahn et al., 2007; Forero-Romero et al., 2009; Libeskind et al., 2017; Paranjape et al., 2018; Ramakrishnan et al., 2019; Alam et al., 2019; Favole et al., 2022; Osato & Okumura, 2023). In particular, tidal anisotropy, which quantifies the degree to which the local gravitational field is directionally dependent, encodes information about the geometry and dynamics of the matter distribution surrounding galaxies and halos. Tidal anisotropy has been shown to correlate strongly with halo assembly bias, concentration, shape, and large-scale bias (Ramakrishnan et al., 2019).
The theoretical motivation for using the tidal field traces back to the nature of initial conditions. Perturbations in Gaussian random fields are inherently triaxial (Zel’dovich, 1970; Doroshkevich, 1970; Bardeen et al., 1986; Heavens & Peacock, 1988; Bond & Myers, 1996; Sheth et al., 2001; Jing & Suto, 2002), and the large-scale shear field plays a central role in shaping the cosmic web from early-time Gaussian fluctuations to the highly nonlinear filamentary structure observed at late times (e.g. Hoffman, 1986; Peebles, 1990; Dubinski, 1992; Bertschinger & Jain, 1994; Bond et al., 1996; Sheth et al., 2001; Desjacques, 2008; Pogosyan et al., 2009; Hidding et al., 2014; Aung & Cohn, 2016; Musso et al., 2018; Shim et al., 2021).
Observationally, galaxy properties such as star formation rate and morphology are known to correlate with their cosmic web environment. For example, galaxies in filaments in the SDSS DR8 main galaxy sample tend to be less star-forming and exhibit earlier-type morphologies than those in more diffuse environments (O’Kane et al., 2024). However, the correlation between large-scale environment and halo properties is subtle. While large-scale bias is known to correlate more strongly with local overdensity measures like the “halo-centric” $\delta_{R}$ than with halo mass (Abbas & Sheth, 2007; Shi & Sheth, 2018; Repp & Szapudi, 2022), such density-based metrics are largely insensitive to the shape of the surrounding matter field—a limitation overcome by measures such as tidal anisotropy.
@@
The paper is organized as follows: we briefly recap on our approach to modelling galactic $\mathrm{H}\rm{\alpha}$ emission in section §2; we describe ways of quantifying the environment in section §3; we present our results for correlations of ELG properties and the environment in section §4.2, and conclude in §5.
2 Simulating Galactic $\mathrm{H}\rm{\alpha}$ emission
@@
The model assumes that the population of the atomic levels are in steady-state and can be computed within the coronal approximation where excitation occurs either from the HI ground state or by recombination of a bare (HII) proton. These approximations are justified by the relatively low densities and by the transparency of the interstellar medium (ISM). As a result, the total $\mathrm{H}\rm{\alpha}$ luminosity of each subhalo can be expressed as a sum over the independent contributions of each $\mathrm{H}\rm{\alpha}$ emission mechanism, i.e.
@@
These contributions to the $\mathrm{H}\rm{\alpha}$ luminosity are computed from all the gas cells bound to the host subhalo. Throughout the paper, we shall refer to ELGs (or galaxies, for simplicity) as any subhalo with resulting luminosity $L_{\mathrm{H}\rm{\alpha}}>10^{39}\ {\rm erg\,s^{-1}}$. Beyond the local physical parameters obtained from the simulations, the key model parameters are:
  • The fraction $\lambda_{h}$ of gas in the hot phase for star forming cells.
  • The characteristic distance $r_{0}$ for absorption of Lyman photons in galaxies. It is simultaneously used as the distance beyond which stars do not contribute to $L_{\mathrm{H}\rm{\alpha}}^{\text{PE}}$ and the distance from the galactic center over which $L_{\mathrm{H}\rm{\alpha}}^{\text{AGN}}$ is suppressed.
@@
In this paper, we adopt the same (constant) parameter values as Rapoport et al. (2025), i.e. $\lambda_{h}=0.1$, $r_{0}=1\ \text{kpc}$ (physical) and $C_{\text{HII}}=10^{40.8}$. Furthermore, we do not include extinction by dust in the source rest frame nor in the Milky-Way in order to work with the intrinsic luminosities of the galaxies. We focus on the $z=1$ snapshot extracted from the TNG300-1 simulation (Nelson et al., 2017; Marinacci et al., 2018; Naiman et al., 2018; Springel et al., 2017; Pillepich et al., 2017), for which the adopted cosmology is $\Omega_{m}=0.3089,\ \Omega_{\Lambda}=0.6911,\ \Omega_{b}=0.0486,\ h=0.6774$ for the matter, dark energy and baryonic energy densities and the Hubble expansion rate (in units of $100\ {\rm km\,s^{-1}\,{\it h}Mpc^{-1}}$), respectively. Dark matter parent halos are identified by a standard friends-of-friends algorithm with a linking length of 0.2 in units of the mean interparticle spacing. Our fiducial simulation is TNG300-1, which has a box size $L=205\ {\rm{\it h}^{-1}Mpc}$ and contains $2500^{3}$ dark matter particles and an equal number of gas particles. We use the higher resolution TNG50-1 (Nelson et al., 2019a; Pillepich et al., 2019), which evolves $2\times 2160^{3}$ dark matter and gas particles in a box of size $L=35\ {\rm{\it h}^{-1}Mpc}$, to assess the impact of numerical resolution on the correlations reported in this work.
3 Quantifying the galactic environment
3.1 Tidal and strain rate tensor
@@
3.2 Anisotropy parameter $\alpha$
@@
Figure 3: Top: Measurements of the scale-dependent bias $b(k)$ of mock $\mathrm{H}\rm{\alpha}$ emitters with $\mathrm{H}\rm{\alpha}$ luminosity $L_{\mathrm{H}\rm{\alpha}}>10^{41}\ {\rm erg\,s^{-1}}$ when split into terciles of $\rm{\alpha_{T}}$. Results are shown separately for central (solid curves) and satellite galaxies (dashed curves). Bottom: Same as the top panel but for $\rm{\alpha_{V}}$.
@@
for mock $\mathrm{H}\rm{\alpha}$ emitters with luminosity $L_{\mathrm{H}\rm{\alpha}}>10^{41}$ and $L_{\mathrm{H}\rm{\alpha}}>10^{42}\ {\rm erg\,s^{-1}}$. $b(k)$ is a measure of the clustering of galaxies relative to the matter distribution. The galaxy-matter cross-spectrum $P_{\rm gm}(k)$ and the matter power spectrum $P_{\rm mm}(k)$ are computed on the $800^{3}$ cubical grid. Results are shown in Fig. 3 for central and satellite galaxies as the solid and dashed curves, respectively. They have been split into terciles of $\alpha_{\rm T}$ and $\alpha_{\rm V}$ of their parent halos as indicated in the figure.
@@
Although the simulation is not large enough to capture the convergence of $b(k)$ to a constant $b_{1}$ on linear scales, $\mathrm{H}\rm{\alpha}$ emitters residing in tidally sheared (high $\alpha$) parent halos are clearly more biased. In general, $b(k)$ of the satellite galaxies is larger than that of central galaxies because the former preferentially reside in massive dark matter halos. The scale dependence of $b(k)$ becomes apparent for $k\gtrsim 0.1\ {\rm{\it h}Mpc^{-1}}$. While it is different for central than for satellite galaxies, our results consistently show that samples with the lowest values of $\alpha_{\rm T}$ or $\alpha_{\rm V}$ exhibit the strongest positive upturn in $b(k)$ – i.e. most scale dependent bias. This is somewhat at odds with the findings of Ramakrishnan & Paranjape (2020) (i.e., halos with larger values of $\alpha$ have larger values of $b_{2}$) unless bias parameters other than $b_{2}$ contribute significantly to the scale-dependence of $b(k)$ for $k\gtrsim 0.1\ {\rm{\it h}Mpc^{-1}}$.
3.3 Cosmic web classification
Halo-centric measurements of $\Psi_{ij}$ and $\Sigma_{ij}$ can be used to classify the environment of dark matter halos and the mock $\mathrm{H}\rm{\alpha}$ emitters they host. Classifications based on $\Psi_{ij}$ and $\Sigma_{ij}$ are known as T-web and V-web respectively (see Hahn et al., 2007; Hoffman et al., 2012; Libeskind et al., 2012; Pfeifer et al., 2022, for detailed discussions). We define $\lambda_{\text{th}}$ to be the eigenvalue threshold above which the local matter distribution is considered as collapsing along the corresponding principal axis. A value of $\lambda_{\text{th}}\sim\mathcal{O}(1)$ usually returns a visual impression of the cosmic web close to that observed in actual data (Hoffman et al., 2012). Since $\Psi_{ij}$ and $\Sigma_{ij}$ differ in the nonlinear regime, we set $\lambda_{\text{th,V}}=0.4H_{0}/(aHf)\approx 0.515$ and $\lambda_{\text{th,T}}=1.5H_{0}/(aHf)\approx 1.932$ for the strain rate and tidal tensor at $z=1$, respectively. The choice of $\lambda_{\text{th,V}}$ is consistent with Forero-Romero et al. (2009); Martizzi et al. (2019); Pfeifer et al. (2022), who advocated $\lambda_{\text{th}}=0.2-0.4$. Our values of $\lambda_{\text{th,T}}$, $\lambda_{\text{th,V}}$ are chosen such that the V- and T-web yield a similar cosmic web classification.
The eigenvalues of each tensor are computed on a grid following the procedure described in §3.1. They are interpolated in configuration space at the halo CoM position, and in $\rm{R_{G}}$-space according to the halo mass $M$. Finally, we assign an environment mark to each halo depending on the number $N_{\lambda}$ of eigenvalues above the threshold $\lambda_{\text{th}}$, with $N_{\lambda}=0$, 1, 2 and 3 corresponding to voids, sheets, filaments and knots respectively.
Figure 4: The comoving differential halo mass function, conditioned on environment. Solid and dashed curves indicate the classification obtained from the V- and T-tensor, respectively. The vertical line indicates, as in Fig. 1, the mass $M\sim 2\times 10^{11}\ {\rm{\it h}}^{-1}M_{\odot}$ below which the smoothing scale $\rm{R_{G}}$ is poorly resolved.
We noted before that large $\alpha_{\rm T}$ are more strongly clustered (c.f. Fig. 3). Clustering strength is usually thought to be strongly correlated with halo mass, so it is important to separate the effects of environment from those of halo mass. To set the stage, Fig. 4 shows that the most massive halos tend to occur in knots or filaments. The abundance of halos with $M\geq 10^{12}\ {\rm{\it h}}^{-1}M_{\odot}$ is lower by more than an order of magnitude in sheets and voids. In addition, notice that, as mass increases, these abundances decrease more steeply for sheets than for knots or filaments. This finding is the same for both the V- and T-web classifications and will be relevant in the next section, when we discuss spatial clustering.
One can use the topological representation of the cosmic web implemented in DisPerSE (Sousbie, 2011) to assign an additional environmental mark. DisPerSE estimates the density field using the discrete halo or galaxy distribution and identifies critical points. Local maxima and minima of the density field define the nodes and voids of the cosmic web, respectively. Furthermore, filaments correspond to the unique integral lines connecting nodes, while sheets are associated with minima along filaments. DisPerSE classifications are publicly available for galaxies in the TNG300-1 simulation, so we use them to assign an environment to each galaxy.
@@ 3.3 Cosmic web classification
Figure 5: Fraction of ELGs in voids, sheets, filaments and knots as a function of $\mathrm{H}\rm{\alpha}$ luminosity, using the same color and linestyle convention of Fig. 4. We additionally show the classification obtained from DisPerSE (dotted), which predicts fewer ELGs in knots and more in sheets. Due to the limited resolution of the grid, ELGs with parent halo mass smaller than $2\times 10^{11}\ {\rm{\it h}}^{-1}M_{\odot}$ are not included in the classification shown here.

### v2

@@ Abstract
Future near-infrared spectroscopic galaxy surveys will target high-redshift emission-line galaxies to test cosmological models. Deriving optimal constraints from emission-line galaxy clustering hinges on a robust understanding of their environmental dependence. Using the TNG300-1 simulation, we explore the correlation between properties of $\mathrm{H}\rm{\alpha}$ emitters and their environment anisotropy rather than traditional density-based measures. Our galactic $\mathrm{H}\rm{\alpha}$ emission model includes contributions from the warm interstellar medium. The environment anisotropy and type are assigned using a halo mass-dependent smoothing scale. We find that most luminous $\mathrm{H}\rm{\alpha}$ emitters ($L_{\mathrm{H}\rm{\alpha}}>10^{42}\ {\rm erg\,s^{-1}}$) reside in filaments and knots. More generally, $\mathrm{H}\rm{\alpha}$ emitters are more biased in strongly anisotropic environments. While correlations with galactic properties are found to be weak, they are statistically significant for host halo masses $M\lesssim 10^{12}\ \ {\rm{\it h}}^{-1}M_{\odot}$. Our analysis motivates further investigation into how environmental anisotropy influences galaxy evolution, and highlights the potential for leveraging these effects in the analyses of upcoming cosmological surveys.
I Introduction
The next generation of galaxy surveys, including Euclid (Euclid Collaboration et al., 2024d), the Dark Energy Spectroscopic Instrument (DESI) (Collaboration et al., 2016), the Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer (SPHEREx) (Doré et al., 2015) and the Nancy Grace Roman Space Telescope (RST) (Spergel et al., 2015) will provide detailed measurements of emission-line galaxy (ELG) clustering. These data will open new avenues for probing key cosmological scenarios, including the nature of dark matter and dark energy, the physics of the early Universe, and the mass of neutrinos. Interpreting the clustering and physical characteristics of ELGs requires a nuanced understanding of the environments in which they reside. Accurate modeling of ELG clustering further hinges on robust priors for galaxy bias parameters, as recent work has shown that the bias of ELGs is sensitive to both selection effects and the details of galaxy formation physics, motivating the use of simulation-calibrated priors and perturbative approaches (e.g., Merson et al., 2019; Barreira et al., 2020; Zhai et al., 2021; Barreira et al., 2021; Yuan et al., 2022; Marinucci et al., 2023; Euclid Collaboration et al., 2024a; Pei et al., 2024; Madar et al., 2024; Reyes-Peraza et al., 2024; Akitsu, 2024; Ivanov et al., 2024; Garcia-Quintero et al., 2024; Khederlarian et al., 2024; Ortega-Martinez et al., 2025; Ivanov, 2025; Sullivan et al., 2025, for recent studies).
@@ I Introduction
Since there is no unique way to define the environment of galaxies, several methods have been developed to characterize it. Common approaches include local density estimators, such as counts-in-cylinders or adaptive smoothing techniques (Dressler, 1980; Cooper et al., 2006) as well as classifications based on group membership or nearest-neighbor distances (e.g., Baldry et al., 2006; Yang et al., 2009). Beyond simple density-based measures (e.g. Hoyle et al., 2005; Park et al., 2007; Blanton and Moustakas, 2009; Tempel et al., 2011), the large-scale tidal field provides a comprehensive framework for characterizing the cosmic web (Shen et al., 2006; Hahn et al., 2007; Forero-Romero et al., 2009; Libeskind et al., 2017; Paranjape et al., 2018; Ramakrishnan et al., 2019; Alam et al., 2019; Favole et al., 2022; Osato and Okumura, 2023). In particular, tidal anisotropy, which quantifies the degree to which the local gravitational field is directionally dependent, encodes information about the geometry and dynamics of the matter distribution surrounding galaxies and halos. Tidal anisotropy has been shown to correlate strongly with halo assembly bias, concentration, shape, and large-scale bias (Ramakrishnan et al., 2019).
The theoretical motivation for using the tidal field traces back to the nature of initial conditions. Perturbations in Gaussian random fields are inherently triaxial (Zel’dovich, 1970; Doroshkevich, 1970; Bardeen et al., 1986; Heavens and Peacock, 1988; Bond and Myers, 1996; Sheth et al., 2001; Jing and Suto, 2002), and the large-scale shear field plays a central role in shaping the cosmic web from early-time Gaussian fluctuations to the highly nonlinear filamentary structure observed at late times (e.g. Hoffman, 1986; Peebles, 1990; Dubinski, 1992; Bertschinger and Jain, 1994; Bond et al., 1996; Sheth et al., 2001; Desjacques, 2008; Pogosyan et al., 2009; Hidding et al., 2014; Aung and Cohn, 2016; Musso et al., 2018; Shim et al., 2021).
Observationally, galaxy properties such as star formation rate and morphology are known to correlate with their cosmic web environment. For example, galaxies in filaments in the SDSS DR8 main galaxy sample tend to be less star-forming and exhibit earlier-type morphologies than those in more diffuse environments (O’Kane et al., 2024). However, the correlation between large-scale environment and halo properties is subtle. While large-scale bias is known to correlate more strongly with local overdensity measures like the “halo-centric” $\delta_{R}$ than with halo mass (Abbas and Sheth, 2007; Shi and Sheth, 2018; Repp and Szapudi, 2022), such density-based metrics are largely insensitive to the shape of the surrounding matter field—a limitation overcome by measures such as tidal anisotropy.
@@
The paper is organized as follows: we briefly recap our model of galactic $\mathrm{H}\rm{\alpha}$ emission in section §II; describe ways of quantifying the environment in section §III; present our results for correlations of ELG properties and the environment in section §IV.2; and conclude in §V.
II Simulating $\mathrm{H}\rm{\alpha}$ emitters
@@
II.1 The $\mathrm{H}\rm{\alpha}$ emission line model
The model assumes that the population of the atomic levels are in steady-state and can be computed within the coronal approximation where excitation occurs either from the HI ground state or by recombination of a bare (HII) proton. These approximations are justified by the relatively low densities and by the transparency of the interstellar medium (ISM). As a result, the total $\mathrm{H}\rm{\alpha}$ luminosity of each galaxy can be expressed as a sum over the independent contributions of each $\mathrm{H}\rm{\alpha}$ emission mechanism, i.e.
@@ II.1 The $\mathrm{H}\rm{\alpha}$ emission line model
These contributions to the $\mathrm{H}\rm{\alpha}$ luminosity are computed from all the gas cells bound to the host subhalo.
Beyond the local physical parameters obtained from the simulations, the key model parameters are:
  • The fraction $\lambda_{h}$ of gas in the hot phase for star-forming cells.
  • The characteristic distance $r_{0}$ for the absorption of Lyman photons in galaxies. It is simultaneously used as the distance beyond which stars do not contribute to $L_{\mathrm{H}\rm{\alpha}}^{\text{PE}}$ and the distance from the galactic center over which $L_{\mathrm{H}\rm{\alpha}}^{\text{AGN}}$ is suppressed.
@@
While we set $\lambda_{h}=0.1$ for all star-forming cells as in Rapoport et al. (2025), we improve our original model in two ways. First, we account for galaxy-to-galaxy variations in the photon mean free path according to the simple prescription
$\lambda_{\text{mfp}}=\min\left(r_{0}\left(\frac{R_{\text{gas}}}{10\ \text{kpc}}\right)^{3}\left(\frac{M_{\text{gas}}}{10^{10}M_{\odot}/h}\right)^{-1}Z^{-1},R_{\text{gas}}\right)\;,$ (2)
where $R_{\text{gas}}$ is the gas half-mass radius, $M_{\text{gas}}$ is the total galactic gas mass, and $Z$ is the mass-weighted average gas metallicity, in solar units.
Second, we explicitly model galactic $\mathrm{H}\rm{\alpha}$ dust extinction using a metallicity-dependent dust cross-section that links the gas content and metallicity of a galaxy to the dust attenuation of its $\mathrm{H}\rm{\alpha}$ emission. Further details will be presented in a forthcoming paper. We fit $r_{0}$ and $C_{\text{HII}}$ to measurements of the $\mathrm{H}\rm{\alpha}$ luminosity functions at $z=1$, obtaining $r_{0}=0.72\ \rm{kpc}$ (physical) and $C_{\text{HII}}=10^{41.36}\ \left[\frac{\ {\rm erg\,s^{-1}}}{M_{\odot}\text{yr}^{-1}}\right]$ . This calibration fully determines the model parameters. Note that the best-fit value of $r_{0}$ implies a cross section (per H-atom) $\sigma\sim 10^{-21}\ \rm{cm^{2}/H}$ consistent with typical dust-absorption cross sections for UV/Lyman-$\alpha$ radiation (Draine, 2011).
II.2 Numerical simulations
We focus on the $z=1$ snapshot extracted from the TNG300-1 simulation (Nelson et al., 2021, 2017; Marinacci et al., 2018; Naiman et al., 2018; Springel et al., 2017; Pillepich et al., 2017), for which the adopted cosmology is $\Omega_{m}=0.3089,\ \Omega_{\Lambda}=0.6911,\ \Omega_{b}=0.0486,\ h=0.6774$ for the matter, dark energy and baryonic energy densities and the Hubble expansion rate (in units of $100\ {\rm km\,s^{-1}\,{\it h}Mpc^{-1}}$), respectively. Dark matter parent halos are identified by a standard friends-of-friends (FoF) algorithm with a linking length of 0.2 in units of the mean interparticle spacing. Our fiducial simulation is TNG300-1, which has a box size $L=205\ {\rm{\it h}^{-1}Mpc}$ and contains $2500^{3}$ dark matter particles and an equal number of gas particles. We use the higher resolution TNG50-1 (Nelson et al., 2019b; Pillepich et al., 2019), which evolves $2\times 2160^{3}$ dark matter and gas particles in a box of size $L=35\ {\rm{\it h}^{-1}Mpc}$, to assess the impact of numerical resolution on the correlations reported in this work. We incorporate our $\mathrm{H}\rm{\alpha}$ emission line model into all the subhalos resolved in the simulations. Throughout the paper, we identify $\mathrm{H}\rm{\alpha}$ emitters as suhalos with resulting luminosity $L_{\mathrm{H}\rm{\alpha}}>10^{39}\ {\rm erg\,s^{-1}}$. For brevity, we will also henceforth refer to them as ELGs.
II.3 Halo Occupation Distribution
The TNG suite of simulations includes information on whether an ELG is a central or a satellite galaxy. This allows us to compute the halo occupation distribution (HOD) of central and satellite ELGs, $N_{c}$ and $N_{s}$, as a function of the host halo mass $M$. The HOD framework provides a fast and flexible approach to modelling galaxy clustering (see Cooray and Sheth, 2002, for a review). For illustration, we show in Fig. 1 the HOD of ELGs identified in the $z=1$ snapshot of TNG300-1. The average central occupation number $\langle N_{c}|M\rangle$ approaches unity at high halo mass and exhibits a local peak at $\sim 10^{12}M_{\odot}/h$. The average satellite occupation number $\langle N_{s}|M\rangle$ increases with $M$, with a shallower rise for rich groups and clusters ($M\gtrsim 10^{14}\ {\rm{\it h}}^{-1}M_{\odot}$). The fraction of halos which host a satellite ELG when the central galaxy is not an ELG is small ($0.01\%$).
Figure 1: The average central and satellite occupation numbers $\langle N_{c}|M\rangle$, $\langle N_{s}|M\rangle$ computed for our $\mathrm{H}\rm{\alpha}$ emission model for TNG300-1 at $z=1$.
III Quantifying the galactic environment
III.1 Tidal and strain rate tensor
@@
III.2 Anisotropy parameter $\alpha$
@@
Figure 4: Top: Measurements of the scale-dependent bias $b(k)$ of mock ELGs with $\mathrm{H}\rm{\alpha}$ luminosity $L_{\mathrm{H}\rm{\alpha}}>10^{41}\ {\rm erg\,s^{-1}}$ when split into terciles of $\rm{\alpha_{T}}$. Results are shown separately for central (solid curves) and satellite galaxies (dashed curves). Bottom: Same as the top panel but for $\rm{\alpha_{V}}$.
@@
for mock ELGs with luminosity $L_{\mathrm{H}\rm{\alpha}}>10^{41}$ and $L_{\mathrm{H}\rm{\alpha}}>10^{42}\ {\rm erg\,s^{-1}}$. $b(k)$ is a measure of the clustering of galaxies relative to the matter distribution. The galaxy-matter cross-spectrum $P_{\rm gm}(k)$ and the matter power spectrum $P_{\rm mm}(k)$ are computed on the $800^{3}$ cubical grid. Results are shown in Fig. 4 for central and satellite galaxies as the solid and dashed curves, respectively. They have been split into terciles of $\alpha_{\rm T}$ and $\alpha_{\rm V}$ of their parent halos as indicated in the figure.
@@
Although the simulation is not large enough to capture the convergence of $b(k)$ to a constant $b_{1}$ on linear scales, ELGs residing in tidally sheared (high $\alpha$) parent halos are clearly more biased. In general, $b(k)$ of the satellite galaxies is larger than that of central galaxies because the former preferentially reside in massive dark matter halos (c.f. Fig. 1). The scale dependence of $b(k)$ becomes apparent for $k\gtrsim 0.1\ {\rm{\it h}Mpc^{-1}}$. While it is different for central than for satellite galaxies, our results consistently show that samples with the lowest values of $\alpha_{\rm T}$ or $\alpha_{\rm V}$ exhibit the strongest positive upturn in $b(k)$ – i.e. most scale dependent bias. This is somewhat at odds with the findings of Ramakrishnan and Paranjape (2020) (i.e., halos with larger values of $\alpha$ have larger values of $b_{2}$) unless bias parameters other than $b_{2}$ contribute significantly to the scale-dependence of $b(k)$ for $k\gtrsim 0.1\ {\rm{\it h}Mpc^{-1}}$.
@@
Figure 5: Left panel: Comoving differential halo mass function, conditioned on environment. Solid and dashed curves indicate the classification obtained from the V- and T-tensors, respectively. Vertical line indicates, as in Fig. 2, the mass $M\sim 2\times 10^{11}\ {\rm{\it h}}^{-1}M_{\odot}$ below which the smoothing scale $\rm{R_{G}}$ is poorly resolved. Right panel: Fraction of ELGs in voids, sheets, filaments and knots as a function of $\mathrm{H}\rm{\alpha}$ luminosity (color and linestyle conventions same as left panel). The classification obtained from DisPerSE (dotted) predicts fewer ELGs in knots and more in sheets. Due to the limited resolution of the grid, ELGs with parent halo mass smaller than $2\times 10^{11}\ {\rm{\it h}}^{-1}M_{\odot}$ are not included here.
III.3 Cosmic web classification
Halo-centric measurements of $\Psi_{ij}$ and $\Sigma_{ij}$ can be used to classify the environment of dark matter halos and the ELGs they host. Classifications based on $\Psi_{ij}$ and $\Sigma_{ij}$ are known as T-web and V-web respectively (see Hahn et al., 2007; Hoffman et al., 2012; Libeskind et al., 2012; Pfeifer et al., 2022, for detailed discussions). We define $\lambda_{\text{th}}$ to be the eigenvalue threshold above which the local matter distribution is considered as collapsing along the corresponding principal axis. A value of $\lambda_{\text{th}}\sim\mathcal{O}(1)$ usually returns a visual impression of the cosmic web close to that observed in actual data (Hoffman et al., 2012). Since $\Psi_{ij}$ and $\Sigma_{ij}$ differ in the nonlinear regime, we set $\lambda_{\text{th,V}}=0.4H_{0}/(aHf)\approx 0.515$ and $\lambda_{\text{th,T}}=1.5H_{0}/(aHf)\approx 1.932$ for the strain rate and tidal tensor at $z=1$, respectively. The choice of $\lambda_{\text{th,V}}$ is consistent with Forero-Romero et al. (2009); Martizzi et al. (2019); Pfeifer et al. (2022), who advocated $\lambda_{\text{th}}=0.2-0.4$. Our values of $\lambda_{\text{th,T}}$, $\lambda_{\text{th,V}}$ are chosen such that the V- and T-web yield a similar cosmic web classification.
The eigenvalues of each tensor are computed on a grid following the procedure described in §III.1. They are interpolated in configuration space at the halo CoM position, and in $\rm{R_{G}}$-space according to the halo mass $M$. Finally, we assign an environment mark to each halo depending on the number $N_{\lambda}$ of eigenvalues above the threshold $\lambda_{\text{th}}$, with $N_{\lambda}=0$, 1, 2 and 3 corresponding to voids, sheets, filaments and knots respectively.
We noted above that halos with large $\alpha_{\rm T}$ or $\alpha_{\rm V}$ are more strongly clustered (c.f. Fig. 4). Clustering strength is usually thought to be strongly correlated with halo mass, so it is important to separate the effects of environment from those of halo mass. To set the stage, the left panel of Fig. 5 shows that the most massive halos tend to occur in knots or filaments. The abundance of halos with $M\geq 10^{12}\ {\rm{\it h}}^{-1}M_{\odot}$ is lower by more than an order of magnitude in sheets and voids. In addition, notice that, as mass increases, these abundances decrease more steeply for sheets than for knots or filaments. This finding is the same for both the V- and T-web classifications and will be relevant in the next section, when we discuss spatial clustering.
One can use the topological representation of the cosmic web implemented in DisPerSE (Sousbie, 2011) to assign an additional environmental mark. DisPerSE estimates the density field using the discrete halo or galaxy distribution and identifies critical points. Local maxima and minima of the density field define the knots and voids of the cosmic web, respectively. Furthermore, filaments correspond to the unique integral lines connecting knots, while sheets are associated with minima along filaments. DisPerSE classifications are publicly available for galaxies in the TNG300-1 simulation, so we use them to assign an environment to each galaxy.